### v1

@@ 4 Discussion
Table 1: Comparisons of the learning capabilities of different local learning methods and their test accuracy [%] on CIFAR-10 and STL-10 dataset.
Method (Self)Unsupervised [^1] Self-supervised or unsupervised Local Sequential [^2] Can handle sequential data or not CIFAR-10 STL-10
SimCLR (Chen et. al. 2020 [49]) 94.0 89.7
PNN-SSL (Laydevant et. al. 2023 [62]) 77.0 -
Bio-SSL (Tang et. al. 2022 [63]) 72.7 68.8
CLAPP (Illing et. al. 2021 [64]) - 73.6
SigProp (Kohan et. al. 2023 [5]) 91.6 -
EqProp (Laborieux et. al. 2021 [65]) 88.6 -
DualProp (Høier et. al. 2023 [66]) 92.3 -
FF (Hinton et. al. 2023 [19]) 59.0 -
FF (Papachristodoulou et. al. 2024 [33]) 78.1 -
FF (Wu et. al. 2024 [44]) 84.7 -
PEPITA (Srinivasan et. al. 2023 [8]) 53.8 -
Act. Learning (Zhou et. al. 2022 [7]) 58.7 -
HardHebb (Miconi et. al. 2021 [11]) 64.8 -
HardHebb (Lagani et. al. 2022 [12]) 64.6 -
Hebb-CHU (Krotov et. al. 2019 [13]) 50.8 -
Hebb-PNorm (Grinberg et. al. 2019 [67]) 72.2 -
SoftHebb (Journé et. al. 2022 [9]) 80.3 76.2
SCFF (ours) 80.8 77.3
@@ 4.1 Comparison to the original FF algorithm
In addressing the limitations of the original FF algorithm, our method introduces several key improvements. Firstly, we have developed an approach for generating negative examples that can be applied to any database. This approach is also biologically plausible, since it operates by aggregating two similar or different images at the input, very much like our eyes do. This innovation directly addresses the criticisms highlighted in[8], which pointed out the biological implausibility of the negative examples used in the original FF algorithm.
@@
Furthermore, we have expanded the applicability of FF to complex unsupervised tasks beyond the MNIST dataset. The SCFF method achieves state-of-the-art (SOTA) accuracy for local methods on challenging datasets such as CIFAR-10 and STL-10, largely outperforming the original FF algorithm (see Table 1). This is a significant advancement, as it demonstrates that our method performs comparably to other local and forward-only algorithms in complex visual tasks, thereby broadening the scope and utility of FF to unlabeled data processing.
@@
Additionally, we have extended the capabilities of FF to include sequential task solving. This extension is crucial for applications in time-series analysis and other domains where data is inherently sequential. By incorporating these improvements, our SCFF method not only overcomes the original limitations of FF but also sets a new benchmark for unsupervised learning algorithms in terms of versatility and performance.
@@ 4.2 Comparison to SOTA self-supervised learning (SSL)
Our SCFF method draws significant inspiration from self-supervised contrastive learning techniques[49, 68]. While the accuracy of SCFF may be lower compared to these methods, primarily due to its layer-wise learning in a purely local manner, it offers unique advantages. Unlike global self-supervised methods, SCFF operates without requiring auxiliary heads (multi-layer nonlinear projector) or complex regularization techniques, which simplifies its implementation and makes it more suitable for neuromorphic computing applications.
@@
Recent developments in local versions of contrastive self-supervised learning have shown promising results [62, 63]. For instance, Laydevant et al. [62] empirically demonstrated that layer-wise SSL objectives can be optimized rather than a single global one, achieving performance comparable to global optimization on datasets such as MNIST and CIFAR-10 (see Table 1). However, the layer-wise training methods involving multi-layer MLP as projector heads might offer better performance in certain tasks, but at the cost of increased computational complexity. Illing et al. [64] have shown that local plasticity rules, when applied through the CLAPP model, can successfully build deep hierarchical representations without the need for backpropagation. However, this method introduces additional processing along the time axis, which may add complexity when dealing with data that lacks temporal dynamics.
4.3 Comparison to other forward-only methods including non-Hebbian and Hebbian based
@@
Recently, other purely forward learning techniques have been developed, driven by their potential for biologically plausible and neuromorphic computing applications [8, 7]. Similar to Forward-Forward (FF), Pepita [8] processes data samples in two forward passes. The first pass is identical to FF, while the input of the second pass is modulated by incorporating information about the error from the first forward pass through top-down feedback. Activation Learning [7] builds on Hebb’s well-known proposal, discovering unsupervised features through local competitions among neurons. However, these methods do not yet scale to more complex tasks, limiting their potential applications.
@@
Recent advances in Hebbian deep learning have also shown remarkable progress [11, 12, 69, 9]. These methods are purely local in space and can be applied purely locally in time, offering a biologically plausible approach to learning. Miconi [11] demonstrated that Hebbian learning in hierarchical convolutional neural networks can be implemented with modern deep learning frameworks by using specific losses whose gradients produce the desired Hebbian updates. However, adding layers has not resulted in significant performance improvements on standard benchmarks [11, 12] (see Table 1). Journe et al. [9] proposed using a simple softmax to implement a soft Winner-Takes-All (WTA) and derived a Hebbian-like plasticity rule (SoftHebb). With techniques like triangle activation and adjustable rectified polynomial units, SoftHebb achieves increased efficiency and biological compatibility, enhancing accuracy compared to state-of-the-art biologically plausible learning methods.
@@ 4.4 Comparison to energy-based learning methods
Energy-based learning methods, such as Equilibrium Propagation (EP), Dual Propagation (DP) and Latent Equilibrium (LE) [70, 66, 71], also offer locality in space and time. These methods have a significant advantage over SCFF due to their strong mathematical foundations, closely approximating gradients from BP and backpropagation through time (BPTT). This theoretical rigor allows them to be applied to a wide range of physical systems, making them particularly appealing for neuromorphic computing applications. EP, for instance, can operate in an unsupervised manner, while recent advancements in Genralized Latent Equilibrium (GLE) [72] have extended these models to handle sequential data effectively.
@@ 6 Methods
The positive and negative examples (e.g. $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ and $\boldsymbol{x}_{j,\text{neg}}^{(0)}$) are generated by concatenating two identical images for the positive examples and two different images for the negative examples. After being processed by the first layer, the output vectors $\boldsymbol{y}_{i,\text{pos}}^{(0)}$ and $\boldsymbol{y}_{j,\text{neg}}^{(0)}$ are obtained. There are two approaches for generating the inputs to the next layer. The first approach is to directly use the first layer’s output of the positive example $\boldsymbol{y}_{i,\text{pos}}^{(0)}$ as the positive input $\boldsymbol{x}_{i,\text{pos}}^{(1)}$, and the first layer output of the negative example $\boldsymbol{y}_{j,\text{neg}}^{(0)}$ as the negative input $\boldsymbol{x}_{j,\text{neg}}^{(1)}$ for the next layer (refer to the highlighted blue section in Algorithm 1 in Appendix C). The second approach involves re-concatenating to form new positive and negative inputs for the next layer. This is done by treating the first layer’s positive outputs as a new dataset and recreating the corresponding positive and negative examples, similar to how the original dataset was processed to generate the initial positive and negative examples (refer to the highlighted blue section in Algorithm 2 in Appendix C).
Appendix C details the workflows of Algorithm 1 and Algorithm 2, focusing on their different approaches to generating positive and negative examples after the first layer. In practice, Algorithm 1 tends to be more effective for training the lower layers immediately following the first layer, while Algorithm 2 shows better performance in training deeper layers. Specifically, for the CIFAR-10 dataset, Algorithm 1 is utilized to train the second layer, while Algorithm 2 is applied to train the third layer. Similarly, for the STL-10 dataset, Algorithm 1 is employed for training the second and third layers, and Algorithm 2 is used for the fourth layer.
@@
Training with the FF loss can lead to excessively high output activations for positive examples, which significantly drives positive gradients and encourages unchecked growth in their activation. To mitigate this, we introduce a small penalty term—the Frobenius Norm of the Goodness vector—into the training loss function. For outputs from a CNN layer, the goodness vector $G_{i,h,w,\text{pos}}^{(l)}$ is a two-dimensional matrix where each element represents the goodness calculated over the channel outputs processed by all filters under the same receptive field. In the case of Bi-RNN outputs, the goodness vector $G_{i,t,\text{pos}}^{(l)}$ is a one-dimensional matrix, with each element representing the goodness at each time step. When a large goodness value is computed for a positive example, it generates a negative gradient that reduces the activation, thereby preventing excessive growth. The impact of this penalty term on training performance is further analyzed in Appendix F.
@@
Appendix D provides detailed information on the specific architecture of this additional pooling layer for various tasks.
@@
Code availability The code to reproduce the results will be made available upon publication.

### v2

@@ 4.1 Comparison to the original FF algorithm
In SCFF, we have expanded the applicability of FF to complex unsupervised tasks beyond the MNIST dataset. The SCFF method achieves state-of-the-art (SOTA) accuracy for local methods on challenging datasets such as CIFAR-10, Tiny ImageNet and STL-10, largely outperforming the original FF algorithm (see Table 1). This is a significant advancement, as it demonstrates that the method performs comparably to other local and forward-only algorithms in complex visual tasks, thereby broadening the scope and utility of FF to unlabeled data processing.
@@
We have also shown that FF can solve sequential tasks. This extension is crucial for applications in time-series analysis and other domains where data is inherently sequential. By incorporating these improvements, our SCFF method not only overcomes the original limitations of FF but also sets a new benchmark for unsupervised learning algorithms in terms of versatility and performance.
4.2 Analysis of the negative examples
The improvements brought by SCFF rely on a novel approach for consistently constructing positive and negative examples, in a way that can be applied to any database.
Figure 5: Probability distributions of relative positions between positive and negative examples. a Theoretical distributions of positive examples from two different classes with distinct means ($02\mu_{1}=$ 0 and $152\mu_{2}=$ 15) and identical variance ($42\Sigma=$ 4) are shown with blue and orange curves, respectively. The theoretical distribution of negative examples derived from the two classes using the formula 4.2 is depicted by the grey curve. b Continuous probability density of LDA applied to the IRIS dataset, displaying contours for positive examples in green, red, and blue, and for negative examples in grey.
The effectiveness of the negative examples in SCFF can be understood through the lens of Noise Contrastive Estimation (NCE). In NCE, a key insight is that “the noise distribution should be close to the data distribution, because otherwise, the classification problem might be too easy and would not require the system to learn much about the structure of the data” [76]. Our method of generating the positive and negative examples aligns with this principle if we treat the negative examples as “noise data”. We assume that the data samples for each class follow a multivariate Gaussian distribution with a shared covariance matrix $\Sigma$ and that each class is statistically independent of the others—assumptions commonly employed in various statistical models [77].
Since the input weight matrices are identical, i.e., $W=W_{1}=W_{2}$, the input transformation simplifies to: $\boldsymbol{y}_{i,\text{pos}}^{(0)}=f(W(\boldsymbol{x}_{k}+\boldsymbol{x}_{k}))=f(W(2\boldsymbol{x}_{k}))$, $\boldsymbol{y}_{j,\text{neg}}^{(0)}=f(W(\boldsymbol{x}_{k}+\boldsymbol{x}_{n})).$ Thus, the positive and negative examples can be interpreted as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{x}_{i,\text{pos}}^{(0)}=2\boldsymbol{x}_{k},\quad\boldsymbol{x}_{j,\text{neg}}^{(0)}=\boldsymbol{x}_{k}+\boldsymbol{x}_{n}.}$ (6)
Therefore, the distributions of positive examples $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ and negative examples $\boldsymbol{x}_{j,\text{neg}}^{(0)}$ follow:
where $\boldsymbol{\mu}_{1}$ and $\boldsymbol{\mu}_{2}$ are means of two different classes respectively.
Theoretically, the negative examples always lie somewhere between two different clusters of positive examples in the sample space, as illustrated in Fig. 5a for the one-dimensional case. Unlike conventional contrastive learning methods that require explicit hard negative sampling, SCFF inherently constructs negative examples that are statistically distinct yet sufficiently similar to the data distribution. This ensures that SCFF maintains a strong contrastive signal without relying on additional augmentation strategies. For practical analysis with a real-world dataset, we visualized the distributions of positive and negative examples from the IRIS dataset [78] using 2D linear discriminant analysis (LDA), which distinguishes between three different types of irises, as shown in Fig.5b. This visualization shows that the negative examples are positioned between different clusters of positive examples, suggesting that they contribute to pushing apart and further separating adjacent positive examples as they are mapped into higher-dimensional space during training. Additionally, negative examples are formed by combining two examples from different classes, enriching the diversity of negative examples and leading to more robust training. For a detailed analysis of how the LDA components evolve during training as the input data is mapped into the feature space and more theoretical results, please refer to Appendix B.
@@ 4.3 Comparison to SOTA self-supervised learning (SSL)
The SCFF method is inspired from self-supervised contrastive learning techniques [55, 79]. While its purely local, layer-wise learning may limit its accuracy compared to end-to-end trained SSL models (Table 2), this design offers unique advantages. SCFF indeed operates without the auxiliary heads (multi-layer nonlinear projector) nor the complex regularization techniques required in global SSL methods, which simplifies its implementation and makes it more suitable for neuromorphic computing applications. SCFF’s local loss and layer normalization also ensure stable activations and mitigate the vanishing gradient problems that typically arise in deep networks trained via backpropagation.
Another key distinction between SCFF and standard contrastive learning approaches lies in the way how negative samples are handled. Traditional methods such as SimCLR [55] and MoCo [79] require large batches or memory banks to construct multiple negative pairs, often incorporating hard negatives and strong augmentations to enhance contrastive separation. These operations are computationally expensive and time-consuming. In contrast, SCFF maintains a fixed 1:1 ratio of positive to negative samples by optimizing a goodness-based objective, eliminating the need for additional transformations or augmentations. This formulation ensures that SCFF remains computationally efficient while still preserving a meaningful contrastive signal. Future work could explore enhancements such as integrating top-down feedback connections and incorporating feedback-informed negative selection, where model predictions dynamically refine the selection of negative samples to enhance contrastive separation to further scale SCFF to modern deep networks like ResNet-50 and Vision Transformers (ViTs).
@@
Recent developments in local versions of contrastive self-supervised learning have shown promising results [80, 81]. For instance, Laydevant et al. [80] empirically demonstrated that layer-wise SSL objectives can be optimized rather than a single global one, achieving performance comparable to global optimization on datasets such as MNIST and CIFAR-10 (see Table 2). However, this accuracy comes from multi-layer MLPs as projector heads at each layer, increasing the computational complexity. Illing et al. [13] have shown that local plasticity rules, when applied through the CLAPP model, can successfully build deep hierarchical representations without the need for backpropagation. However, this method introduces additional processing along the time axis, which may add complexity when dealing with data that lacks temporal dynamics.
Table 2: Test accuracy [%] (top-1) comparison of SCFF with self-supervised Learning methods on MNIST, CIFAR-10, STL-10, and Tiny ImageNet datasets. The symbols , , and - mean “yes”, “no”, and “no reported results” respectively.
Method Unsuper -vised MNIST CIAFR -10 STL -10 Tiny -ImageNet
SimCLR (Chen et. al. 2020 [55]) - 94.0 89.7 53.4 [82]
Bio-SSL (Tang et. al. 2022 [81]) - 72.7 68.8 -
PNN-SSL (Laydevant et. al. 2023 [80]) 96.6 77.0 - -
CLAPP (Illing et. al. 2021 [13]) - - 73.6 -
SCFF (ours) 98.7 80.8 77.3 35.7
4.4 Comparison to other forward-only methods
The development of purely forward learning techniques has been historically driven by their potential for biologically plausible and neuromorphic computing applications [60, 61]. We compare recent and SOTA results of purely-forward methods with local learning rules to SCFF in Table 3.
@@ 4.4 Comparison to other forward-only methods
Similar to Forward-Forward (FF), Pepita [60] processes data samples in two forward passes. The first pass is identical to FF, while the input of the second pass is modulated by incorporating information about the error from the first forward pass through top-down feedback. Activation Learning [61] builds on Hebb’s well-known proposal, discovering unsupervised features through local competitions among neurons. However, these methods do not yet scale to more complex tasks, limiting their potential applications.
Table 3: Test accuracy [%] comparison of SCFF with other forward-only methods on MNIST (top-1), CIFAR-10 (top-1), STL-10 (top-1) datasets, and Tiny ImageNet (top-5). The symbols , , and - mean “yes”, “no”, and “no reported results” respectively.
Method Unsuper -vised MNIST CIAFR -10 STL -10 Tiny ImageNet
SigProp (Kohan et. al. 2023 [50]) 98.2 91.6 - -
PEPITA (Srinivasan et. al. 2023 [60]) 98.4 53.8 - -
Act. Learning (Zhou et. al. 2022 [61]) 97.1 58.7 - -
HardHebb (Miconi et. al. 2021 [57]) - 64.8 - -
HardHebb (Lagani et. al. 2021 [12]) 98.5 [83] 65.9 - 37.0
Hebb-CHU (Krotov et. al. 2019 [84]) 98.5 50.8 - -
Hebb-PNorm (Grinberg et. al. 2019 [85]) - 72.2 - -
SoftHebb (Journé et. al. 2022 [11]) 97.8 80.3 76.2 -
SCFF (ours) 98.7 80.8 77.3 59.8
Table 4: Single-layer training performance (in terms of epoch duration) for different publicly available forward-only and local learning repositories, compared to SCFF, on CIFAR-10. Accuracy results of the overall models from the respective papers on CIFAR-10 are also reported. All experiments were conducted on a single NVIDIA RTX 4090 GPU.
Method Unsupervised Epoch duration (s) Acc. (%)
SigProp (Kohan et. al. 2023 [50]) 7.5 91.6
HardHebb (Miconi et. al. 2021 [57]) 1.6 64.8
HardHebb (Lagani et. al. 2022 [83]) 5.1 64.6
SoftHebb (Journé et. al. 2022 [11]) 2.6 80.3
SCFF (ours) 1.2 80.8
@@
Hebbian deep learning has also achieved remarkable progress recently [57, 83, 86, 11]. These methods are purely local in space and can be applied purely locally in time, offering a biologically plausible approach to learning. Miconi [57] demonstrated that Hebbian learning in hierarchical convolutional neural networks can be implemented with modern deep learning frameworks by using specific losses whose gradients produce the desired Hebbian updates. However, adding layers has not resulted in significant performance improvements on standard benchmarks [57, 83]. Journe et al. [11] proposed using a simple softmax to implement a soft Winner-Takes-All (WTA) and derived a Hebbian-like plasticity rule (SoftHebb). With techniques like triangle activation and adjustable rectified polynomial units, SoftHebb achieves increased efficiency and biological compatibility, enhancing accuracy compared to state-of-the-art biologically plausible learning methods.
@@
Another notable distinction between SCFF and SoftHebb lies in the way the classifier is applied to the neural network. SoftHebb attaches a linear classifier to the final layer, utilizing all neurons from the last layer as inputs. In contrast, SCFF applies an additional pooling operation to aggregate information from intermediate layers, reducing the total number of input neurons before feeding them into the linear classifier. This results in a more compact and efficient representation. For example, on CIFAR-10, SoftHebb employs 24,576 input neurons for classification, whereas SCFF only uses 18,432 due to the added pooling operation. Similarly, for STL-10, SoftHebb utilizes 221,184 input neurons, while SCFF significantly reduces this number to 38,400. This reduction in the number of neurons helps simplify the classification process and contributes to computational efficiency.
Additionally, our method provides a computational advantage. Table 4 presents a comparison of single-layer training time per epoch on the CIFAR-10 dataset. To ensure a fair comparison, all methods are evaluated under a common setting: a single convolutional layer with $5\times 5$ filters, 3 input channels, and 96 output channels, using their respective training methodologies. The results show that SCFF enables more efficient training of convolutional layers compared to prior methods. This efficiency stems from SCFF’s learning rule, which eliminates weight transpositions and matrix multiplications during gradient computation, thereby reducing computational complexity. In contrast, algorithms such as SoftHebb rely on these operations, leading to increased training time.
@@ 4.5 Comparison to energy-based learning methods
Energy-based learning methods (Table 5), such as Equilibrium Propagation (EP), Dual Propagation (DP) and Latent Equilibrium (LE) [30, 10, 87], also offer locality in space and time. These methods have a significant advantage over SCFF due to their strong mathematical foundations, closely approximating gradients from BP and backpropagation through time (BPTT). This theoretical rigor allows them to be applied to a wide range of physical systems, making them particularly appealing for neuromorphic computing applications. EP, for instance, can operate in an unsupervised manner [88], while recent advancements in Genralized Latent Equilibrium (GLE) [89] have extended these models to handle sequential data effectively.
@@
Table 5: Test accuracy [%] comparison of SCFF with energy-based methods on MNIST, CIFAR-10 and STL-10 datasets. The symbols , , and - mean “yes”, “no”, and “no reported results” respectively.
Method Unsupervised MNIST CIFAR-10 STL-10
EqProp (Laborieux et. al. 2021 [33]) 98.0 88.6 -
EqProp (Liu et. al. 2024 [88]) 97.6 71.5 -
DualProp (Høier et. al. 2023 [10]) 98.4 92.3 -
SCFF (ours) 98.7 80.8 77.3
@@ 6 Methods
The positive and negative examples (e.g. $\boldsymbol{x}_{i,\text{pos}}^{(0)}$ and $\boldsymbol{x}_{j,\text{neg}}^{(0)}$) are generated by concatenating two identical images for the positive examples and two different images for the negative examples. After being processed by the first layer, the output vectors $\boldsymbol{y}_{i,\text{pos}}^{(0)}$ and $\boldsymbol{y}_{j,\text{neg}}^{(0)}$ are obtained. There are two approaches for generating the inputs to the next layer. The first approach is to directly use the first layer’s output of the positive example $\boldsymbol{y}_{i,\text{pos}}^{(0)}$ as the positive input $\boldsymbol{x}_{i,\text{pos}}^{(1)}$, and the first layer output of the negative example $\boldsymbol{y}_{j,\text{neg}}^{(0)}$ as the negative input $\boldsymbol{x}_{j,\text{neg}}^{(1)}$ for the next layer (refer to the highlighted blue section in Algorithm 1 in Appendix D). The second approach involves re-concatenating to form new positive and negative inputs for the next layer. This is done by treating the first layer’s positive outputs as a new dataset and recreating the corresponding positive and negative examples, similar to how the original dataset was processed to generate the initial positive and negative examples (refer to the highlighted blue section in Algorithm 2 in Appendix D).
Appendix D details the workflows of Algorithm 1 and Algorithm 2, focusing on their different approaches to generating positive and negative examples after the first layer. In practice, Algorithm 1 tends to be more effective for training the lower layers immediately following the first layer, while Algorithm 2 shows better performance in training deeper layers. Specifically, for the CIFAR-10 dataset, Algorithm 1 is utilized to train the second layer, while Algorithm 2 is applied to train the third layer. Similarly, for the STL-10 dataset, Algorithm 1 is employed for training the second and third layers, and Algorithm 2 is used for the fourth layer.
@@
Training with the FF loss can lead to excessively high output activations for positive examples, which significantly drives positive gradients and encourages unchecked growth in their activation. To mitigate this, we introduce a small penalty term—the Frobenius Norm of the Goodness vector—into the training loss function. For outputs from a CNN layer, the goodness vector $G_{i,h,w,\text{pos}}^{(l)}$ is a two-dimensional matrix where each element represents the goodness calculated over the channel outputs processed by all filters under the same receptive field. In the case of Bi-RNN outputs, the goodness vector $G_{i,t,\text{pos}}^{(l)}$ is a one-dimensional matrix, with each element representing the goodness at each time step. When a large goodness value is computed for a positive example, it generates a negative gradient that reduces the activation, thereby preventing excessive growth. The impact of this penalty term on training performance is further analyzed in Appendix G.
@@
Appendix E provides detailed information on the specific architecture of this additional pooling layer for various tasks.
Training setup
All experiments were conducted on a server equipped with an NVIDIA GeForce RTX 4090 GPU (24 GB memory). Training and evaluation were implemented using PyTorch and executed on a single GPU.
@@
Code availability The code to reproduce the results is available at: https://github.com/neurophysics-cnrsthales/contrastive-forward-forward